### v2

Abstract
In this work we propose and analyse a structure-preserving approximation of the non-isothermal Cahn-Hilliard-Navier-Stokes system using conforming finite elements in space and implicit time discretisation with convex-concave splitting. The system is first reformulated into a variational form which reveal the structure of the equations, which is then used in the subsequent approximation.
1 Introduction
The non-isothermal Cahn-Hilliard-Navier-Stokes (CHNST) system has increasingly gained attention for the investigation of different phenomena ranging from two-phase flows to fluid-phase-coupled interactions both of which have relevant importance in scientific and industrial applications, such as additive manufacturing and inkjet printing van2017binary ; Yang2020 ; dadvand2021advected . For instance, the modeling and simulation of powder bed fusion additive manufacturing (PBF-AM) process employs the non-isothermal CHNST to depict coupled processes in PBF-AM such as fluid-phase interaction, melt flow dynamics and heat transfer Yang2020 . For the CHNST system we consider the following system of partial differential equations
$\partial_{t}\phi$ $+\mathbf{u}\cdot\nabla\phi-\operatorname{div}(\mathbf{L}_{11}\nabla\mu+\mathbf{L}_{12}\nabla\theta)=0,\qquad\mu=-\gamma\Delta\phi+\partial_{\phi}\Psi(\phi,\theta),$ (1)
$\partial_{t}e$ $+\mathbf{u}\cdot\nabla e-\operatorname{div}(\mathbf{L}_{12}\nabla\mu-\mathbf{L}_{22}\nabla\theta)-(\eta\mathrm{D}\mathbf{u}-\boldsymbol{\sigma}):\nabla\mathbf{u}=0,$ (2)
$\partial_{t}\mathbf{u}$ $+(\mathbf{u}\cdot\nabla)\mathbf{u}-\operatorname{div}(\eta\mathrm{D}\mathbf{u}-p\mathbf{I}-\boldsymbol{\sigma})=0,\qquad\operatorname{div}(\mathbf{u})=0.$ (3)
The above system is complemented by periodic boundary conditions and initial conditions. In this context $\phi$ denotes a conserved phase-field, $\mathbf{u}$ is the flow velocity, $\theta$ is the inverse temperature and $e\equiv e(\phi,\theta)$ is the internal energy. To close the system we consider the Helmholtz free energy density and the Korteweg stress given by
$\tilde{\Psi}(\phi,\theta)=\Psi(\phi,\theta)+\tfrac{\gamma}{2}|\nabla\phi|^{2},\qquad\boldsymbol{\sigma}:=\tfrac{\gamma}{\theta}\nabla\phi\otimes\nabla\phi.$
From this, one can compute the internal energy and entropy according to Alt1990 ; Alt1992 , i.e. $e=\partial_{\theta}\tilde{\Psi}$ and $\hat{s}(e(\theta,\phi),\phi)=s(\theta,\phi)=\theta e-\tilde{\Psi}$.
The above system is already formulated in the inverse temperature $\theta$ which is more convenient for a finite element discretisation. Transformation from the original temperature to the inverse temperature can be found by Pawlow and Alt, see Alt1992 , and also brunk2023variational for a more complicated non-isothermal Cahn-Hilliard-Allen-Cahn system. The general derivation of non-isothermal phase-field models from thermodynamic sound principles is more delicate and we refer to Charach1998 ; Fabrizio2006 ; Pawlow2016 for several different modelling approaches.
For such kind of systems there are works Guo_2015 ; Sun2020 , where they consider finite differences in space together with an Energy Quadratisation (EQ) ansatz. This renders the driving functional quadratic, which facilities standard time discretisation methods. However, the price of such an ansatz is the relaxation of entropy, which in some cases results in a drift-off from the original entropy. The isothermal version of the above system, i.e. the Cahn-Hilliard-Navier-Stokes system, is typically called model H and was introduced by Hohenberg and Halperin Hohenberg . This model is employed in the context of two-phase flow especially in phase separation. In this case there are many well-established techniques to solve the system with finite elements Diegel2017 ; Feng06 ; Han2015 and finite differences Gong2018 ; ChenZhao2020 ; LiShen2022 in space.
In this work we will provide a corresponding fully discrete method for the Cahn-Hilliard-Navier-Stokes system using standard finite elements in space and suitable explicit and implicit time-stepping. The discretisation can be seen as an natural extension of the well-known convex-concave splitting for the isothermal Cahn-Hilliard-Navier-Stokes system to the non-isothermal case.
The work is structured as follows. In Section 2 we will introduce relevant notation, and derive a variational formulation of System (1)–(3) which is suitable for a finite element approximation. Finally we present the fully discrete method and present the main result, i.e. preservation of total energy conservation and entropy production. Section 3 illustrates the theoretical scheme with a suitable convergence test. In Section 4 we conclude the work and present an outlook for future work.
2 Notation And Main Result
Before we present out discretization method and main results in detail, let us briefly introduce our notation and main assumptions, and recall some basic facts.
[…]
$\langle u,v\rangle=\int_{\Omega}u\cdot v\quad\forall u,v\in L^{2}(\Omega)\quad\text{ with norm }\|u\|_{0}^{2}:=\int_{\Omega}u^{2}.$
We introduce the skew-symmetric formulation of $\mathbf{c}(\mathbf{u},\mathbf{v},\mathbf{w}):=\langle(\mathbf{u}\cdot\nabla)\mathbf{v},\mathbf{w}\rangle$ via
$\mathbf{c}_{skw}(\mathbf{u},\mathbf{v},\mathbf{w})=\tfrac{1}{2}\mathbf{c}(\mathbf{u},\mathbf{v},\mathbf{w})-\tfrac{1}{2}\mathbf{c}(\mathbf{u},\mathbf{w},\mathbf{v})$
with the relevant property that $\mathbf{c}_{skw}(\mathbf{u},\mathbf{v},\mathbf{v})=0$ even if $\mathbf{u}$ is not divergence-free.
We require the following assumptions for this work.
  • (A1) The interface parameters $\gamma$ is a positive constant.
  • (A2) The viscosity function $\eta\equiv\eta(\phi,\theta)$ is strictly positive.
  • (A3) The diffusion matrix $\mathbf{L}\equiv\mathbf{L}(\rho,\nabla\rho,\theta)\in\mathbb{R}^{2d\times 2d}$ given by $\begin{pmatrix}\mathbf{L}_{11}&-\mathbf{L}_{12}\\ -\mathbf{L}_{12}&\mathbf{L}_{22}\end{pmatrix}$ is symmetric and strictly positive definite.
  • (A4) For the driving potential $\Psi(\cdot,\cdot):\mathbb{R}\times\mathbb{R}_{+}\to\mathbb{R}$ we assume that for every fixed $\phi$ the potential $\Psi(\phi,\cdot):\mathbb{R}_{+}\to\mathbb{R}$ is concave and goes to infinity for $\theta\to 0$. Furthermore, for every fixed $\theta\in\mathbb{R}_{+}$ the potential $\Psi(\cdot,\theta)$ can be decomposed in a strictly convex and a strictly concave part, denoted by $\Psi_{vex},\Psi_{cav}.$
Variational Formulation: In this paragraph we will recast the system into a variational form which is directly usable for a conforming finite element discretisation. We will see that the correct formulation of the Korteweg stress $\boldsymbol{\sigma}$ will be crucial. In the isothermal case it is well-known that a reformulation of $\boldsymbol{\sigma}$ and the pressure $p$ in terms of $\phi,\mu,\pi$, where $\pi$ is a modified pressure, allows to deduce such variational formulations and we will provide an non-isothermal expansion which reveal the strong coupling of the temperature.
Lemma 1
A sufficiently regular solution $(\phi,e,\mathbf{u},p)$ of the system (1)–(3) fulfills the following variational formulation
$\langle\partial_{t}\phi,\psi\rangle$ $-\langle\phi\mathbf{u},\nabla\psi\rangle+\langle\mathbf{L}_{11}\nabla\mu-\mathbf{L}_{12}\nabla\theta,\nabla\psi\rangle=0,$
$\langle\mu,\xi\rangle$ $-\gamma\langle\nabla\phi,\nabla\xi\rangle-\langle\partial_{\phi}\Psi,\xi\rangle=0,$
$\langle\partial_{t}e,w\rangle$ $+\langle\mathbf{L}_{12}\nabla\mu-\mathbf{L}_{22}\nabla\theta,\nabla w\rangle-\langle\eta|\mathrm{D}\mathbf{u}|^{2},w\rangle$
$-\langle\boldsymbol{\sigma}\mathbf{u},\nabla w\rangle-\langle\tfrac{\phi}{\theta}\nabla\mu-\boldsymbol{\sigma}\tfrac{\nabla\theta}{\theta},\mathbf{u}w\rangle-\langle s+\phi\mu,\mathbf{u}\cdot\nabla\tfrac{w}{\theta}\rangle=0,$
$\langle\partial_{t}\mathbf{u},\mathbf{v}\rangle$ $+\mathbf{c}_{skw}(\mathbf{u},\mathbf{u},\mathbf{v})+\langle\eta\mathrm{D}\mathbf{u},\mathrm{D}\mathbf{v}\rangle-\langle\pi,\operatorname{div}(\mathbf{v})\rangle$
$+\langle\tfrac{\phi}{\theta}\nabla\mu+(s+\phi\mu)\nabla\tfrac{1}{\theta}-\boldsymbol{\sigma}\tfrac{\nabla\theta}{\theta},\mathbf{v}\rangle=0,$
00 0 $=\langle\operatorname{div}(\mathbf{u}),q\rangle$
for sufficiently regular test functions $\psi,\xi,\mathbf{v},w,q$ and $\pi:=p+e-\frac{s+\phi\mu}{\theta}$.
Proof
The main part of the procedure is standard, i.e. multiplication by test functions, integration by parts. For the Korteweg stress we expand as follows
$\operatorname{div}(\boldsymbol{\sigma})=\operatorname{div}(\tfrac{\gamma}{\theta}\nabla\phi\otimes\nabla\phi)$ $=\tfrac{\gamma}{\theta}(\Delta\phi\nabla\phi+\nabla\tfrac{1}{2}|\nabla\phi|^{2})+\nabla\phi\otimes\nabla\phi\cdot\nabla\tfrac{\gamma}{\theta}$
$=-\tfrac{\mu}{\theta}\nabla\phi+\theta\boldsymbol{\sigma}\nabla\tfrac{1}{\theta}+\tfrac{1}{\theta}(\nabla\tfrac{\gamma}{2}|\nabla\phi|^{2}+\tfrac{\partial\Psi}{\partial\phi}\nabla\phi),$
$=-\tfrac{\mu}{\theta}\nabla\phi+\theta\boldsymbol{\sigma}\nabla\tfrac{1}{\theta}+\tfrac{1}{\theta}(\nabla\tfrac{\gamma}{2}|\nabla\phi|^{2}+\nabla\Psi-e\nabla\theta)$
$=-\tfrac{\mu}{\theta}\nabla\phi+\theta\boldsymbol{\sigma}\nabla\tfrac{1}{\theta}-\tfrac{1}{\theta}\nabla s+\tfrac{1}{\theta}\theta\nabla e$
$=-\tfrac{\mu}{\theta}\nabla\phi+(\theta\boldsymbol{\sigma}+s)\nabla\tfrac{1}{\theta}-\nabla\tfrac{s}{\theta}+\nabla e$
$=\tfrac{\phi}{\theta}\nabla\mu+\nabla\tfrac{1}{\theta}(s+\phi\mu+\theta\boldsymbol{\sigma})-\nabla(\tfrac{s+\phi\mu}{\theta})+\nabla e.$
The Korteweg stress appears in the internal energy equation as well as the Navier-Stokes equation. In weak form we find
$\langle\boldsymbol{\sigma}:\nabla\mathbf{u},w\rangle$ $=-\langle\boldsymbol{\sigma}\mathbf{u},\nabla w\rangle-\langle\operatorname{div}(\boldsymbol{\sigma})\mathbf{u},w\rangle$
$=-\langle\boldsymbol{\sigma}\mathbf{u},\nabla w\rangle-\langle\tfrac{\phi}{\theta}\nabla\mu-\boldsymbol{\sigma}\tfrac{\nabla\theta}{\theta},\mathbf{u}w\rangle-\langle s+\phi\mu,\mathbf{u}\cdot\nabla\tfrac{w}{\theta}\rangle-\langle\mathbf{u}\cdot\nabla e,w\rangle,$
$\langle\operatorname{div}(\boldsymbol{\sigma}),\mathbf{v}\rangle$ $=\langle\tfrac{\phi}{\theta}\nabla\mu+\nabla\tfrac{1}{\theta}(s+\phi\mu)-\boldsymbol{\sigma}\tfrac{\nabla\theta}{\theta},\mathbf{v}\rangle-\langle e-\tfrac{s+\phi\mu}{\theta},\operatorname{div}(\mathbf{v})\rangle.$
Note that after insertion the convective term $\langle\mathbf{u}\cdot\nabla e,w\rangle$ cancels out.
This variational formulation allows to deduce the thermodynamics quantities immediately by inserting suitable test function.
Theorem 2.1
For a sufficiently regular solution $(\phi,\mu,\theta,\mathbf{u},\pi)$ of (1) – (3) conservation of mass and total energy as well as entropy production holds, i.e.
$\langle\partial_{t}\phi,1\rangle$ $=0,\qquad\langle\partial_{t}(\tfrac{1}{2}|\mathbf{u}|^{2}+e(\phi,\theta)),1\rangle=0,$
$\langle\partial_{t}s(\phi,\theta),1\rangle$ $=\|\sqrt{\eta\theta}\mathrm{D}\mathbf{u}\|_{0}^{2}+\langle(\nabla\mu,\nabla\theta)^{\top},\mathbf{L}(\nabla\mu,\nabla\theta)^{\top}\rangle\geq 0.$
Proof
Conservation of mass follows immediately by inserting $\psi=1$. Conservation of total energy is obtain by inserting $w=1$, $\mathbf{v}=\mathbf{u}$ $q=\pi$, using the skew-symmetry $\mathbf{c}_{skw}(\mathbf{u},\mathbf{u},\mathbf{u})=0$, i.e.
$\tfrac{d}{dt}$ $\langle\tfrac{1}{2}|u|^{2}+e,1\rangle=\langle\partial_{t}\mathbf{u},\mathbf{u}\rangle+\langle\partial_{t}e,1\rangle$
$=-\mathbf{c}_{skw}(\mathbf{u},\mathbf{u},\mathbf{u})-\langle\eta\mathrm{D}\mathbf{u},\mathrm{D}\mathbf{u}\rangle+\langle\pi,\operatorname{div}(\mathbf{u})\rangle$
$\quad-\langle\tfrac{\phi}{\theta}\nabla\mu+(s+\phi\mu)\nabla\tfrac{1}{\theta}-\boldsymbol{\sigma}\tfrac{\nabla\theta}{\theta},\mathbf{u}\rangle$
$\quad+\langle\mathbf{L}_{12}\nabla\mu-\mathbf{L}_{22}\nabla\theta,\nabla 1\rangle+\langle\eta|\mathrm{D}\mathbf{u}|^{2},1\rangle+\langle\boldsymbol{\sigma}\mathbf{u},\nabla 1\rangle$
$\quad+\langle\tfrac{\phi}{\theta}\nabla\mu+\boldsymbol{\sigma}\tfrac{\nabla\theta}{\theta},\mathbf{u}\rangle+\langle s+\phi\mu,\mathbf{u}\cdot\nabla\tfrac{1}{\theta}\rangle$
$=0.$
For the entropy production we insert $\psi=-\mu,\xi=-\partial_{t}\phi$ and $w=\theta$ and by cancellation we immediately find
$\langle\partial_{t}s,1\rangle$ $=\langle\partial_{t}e,\theta\rangle-\langle\partial_{t}\phi,\partial_{\phi}\Psi\rangle-\gamma\langle\nabla\phi,\nabla\partial_{t}\phi\rangle=\langle\partial_{t}e,\theta\rangle-\langle\mu,\partial_{t}\phi\rangle$
$=-\langle\phi\mathbf{u},\nabla\mu\rangle+\langle\mathbf{L}_{11}\nabla\mu-\mathbf{L}_{12}\nabla\theta,\nabla\mu\rangle-\langle\mathbf{L}_{12}\nabla\mu+\mathbf{L}_{22}\nabla\theta,\nabla\theta\rangle$
$+\langle\eta|\mathrm{D}\mathbf{u}|^{2},\theta\rangle+\langle\boldsymbol{\sigma}\mathbf{u},\nabla\theta\rangle+\langle\tfrac{\phi}{\theta}\nabla\mu-\boldsymbol{\sigma}\tfrac{\nabla\theta}{\theta},\mathbf{u}\theta\rangle+\langle s+\phi\mu,\mathbf{u}\cdot\nabla 1\rangle$
$=\langle\mathbf{L}_{11}\nabla\mu-\mathbf{L}_{12}\nabla\theta,\nabla\mu\rangle-\langle\mathbf{L}_{12}\nabla\mu-\mathbf{L}_{22}\nabla\theta,\nabla\theta\rangle+\langle\eta|\mathrm{D}\mathbf{u}|^{2},\theta\rangle+$
$=\|\sqrt{\eta\theta}\mathrm{D}\mathbf{u}\|_{0}^{2}+\langle(\nabla\mu,\nabla\theta)^{\top},\mathbf{L}(\nabla\mu,\nabla\theta)^{\top}\rangle=:\mathcal{D}_{\theta,\mathbf{L}}(\mathrm{D}\mathbf{u},\nabla\mu,\nabla\theta)\geq 0.$
Time Discretization: We partition the time interval $[0,T]$ into uniformly with step size $\tau>0$ and introduce $\mathcal{I}_{\tau}:=\{t^{0}=0,t^{1}=\tau,\ldots,t^{n_{T}}=T\}$, where $n_{T}=\tfrac{T}{\tau}$ is the absolute number of time steps. We denote by $\Pi^{1}_{c}(\mathcal{I}_{\tau};X),\Pi^{0}(\mathcal{I}_{\tau};X)$ the spaces of continuous piece-wise linear and piece-wise constant functions on $\mathcal{I}_{\tau}$ with values in the space or set $X$. By $g^{n+1},g^{n},g^{n+1/2}$ we denote the new, the old time level and the midpoint approximation of $g$, i.e. $(g^{n+1}+g^{n})/2$. We introduce the time difference and the discrete time derivative via
$d^{n+1}g=g^{n+1}-g^{n},\qquad d^{n+1}_{\tau}g=\tau^{-1}(g^{n+1}-g^{n})=\tau^{-1}d^{n+1}g.$
Space Discretization: For the spacial discretisation we require that $\mathcal{T}_{h}$ is a geometrically conforming partition of $\Omega$ into simplices that can be extended periodically to periodic extensions of $\Omega$. We denote the space of continuous, piecewise linear and quadratic functions over $\mathcal{T}_{h}$ as well as the space of mean free and set of positive piecewise-linear functions over $\mathcal{T}_{h}$ via
$\mathcal{V}_{h}$ :=$:=\{v\in H^{1}(\Omega)\cap C^{0}(\bar{\Omega}):v|_{K}\in P_{1}(K)\quad\forall K\in\mathcal{T}_{h}\},$
$\mathcal{X}_{h}^{d}$ :=$:=\{v\in H^{1}(\Omega)^{d}\cap C^{0}(\bar{\Omega})^{d}:v|_{K}\in P_{2}(K)^{d}\quad\forall K\in\mathcal{T}_{h}\},$
$\mathcal{Q}_{h}$ :=$:=\{v\in\mathcal{V}_{h}:\langle v,1\rangle=0\},\qquad\mathcal{V}_{h}^{+}:=\{v\in\mathcal{V}_{h}:v(x)>0,\;\forall x\in\Omega\}.$
We introduce the abbreviation for the convex-concave splitting by
$\Psi(\phi_{h}^{n+1},\phi_{h}^{n},\theta_{h}^{n+1})=\Psi_{vex}(\phi_{h}^{n+1},\theta_{h}^{n+1})+\Psi_{cav}(\phi_{h}^{n},\theta_{h}^{n+1})$
and we will abbreviate when suitable $e(\phi_{h}^{n+1},\theta_{h}^{n+1})=:e_{h}^{n+1}$ and similarly for $s,\Psi$.
We then propose the fully discrete time-stepping method for the CHNST system.
Problem 1
Let $(\phi_{h,0},\mathbf{u}_{h,0},\theta_{h,0})\in\mathcal{V}_{h}\times\mathcal{X}_{h}^{d}\times\mathcal{V}_{h}^{+}$ be given. Find the functions $(\phi_{h},\mathbf{u}_{h},\theta_{h})\in\Pi^{1}_{c}(\mathcal{I}_{\tau};\mathcal{V}_{h}\times\mathcal{X}_{h}^{d}\times\mathcal{V}_{h}^{+})$ and $(\mu_{h},\pi_{h})\in\Pi^{0}(\mathcal{I}_{\tau};\mathcal{V}_{h}\times\mathcal{Q}_{h})$ such that
$\langle d_{\tau}^{n+1}\phi_{h},\psi_{h}\rangle$ $-\langle\phi_{h}^{*}\mathbf{u}_{h}^{n+1/2},\nabla\psi_{h}\rangle+\langle\mathbf{L}_{11}^{*}\nabla\mu_{h}^{n+1}-\mathbf{L}_{12}^{*}\nabla\theta_{h}^{n+1},\nabla\psi_{h}\rangle=0,$
$\langle\mu_{h}^{n+1},\xi_{h}\rangle$ $-\gamma\langle\nabla\phi_{h}^{n+1},\nabla\xi_{h}\rangle-\langle\partial_{\phi}\Psi(\phi^{n+1}_{h},\phi^{n}_{h},\theta^{n+1}_{h}),\xi_{h}\rangle=0,$
$\langle d_{\tau}^{n+1}e_{h},w_{h}\rangle$ $-\langle\eta^{*}|\mathrm{D}\mathbf{u}_{h}^{n+1/2}|^{2},w_{h}\rangle+\langle\mathbf{L}_{12}^{*}\nabla\mu_{h}^{n+1}-\mathbf{L}_{22}^{*}\nabla\theta_{h}^{n+1},\nabla w_{h}\rangle$
$-\langle\boldsymbol{\sigma}_{h}^{*}\cdot\mathbf{u}_{h}^{n+1/2},\nabla w_{h}\rangle-\langle\tfrac{\phi_{h}^{*}}{\theta_{h}^{n+1}}\nabla\mu_{h}^{n+1}-\boldsymbol{\sigma}_{h}^{*}\tfrac{\nabla\theta_{h}^{n+1}}{\theta_{h}^{n+1}},\mathbf{u}_{h}^{n+1/2}w_{h}\rangle$
$-\langle(s_{h}^{*}+\phi_{h}^{*}\mu_{h}^{*})\mathbf{u}_{h}^{n+1/2},\tfrac{\theta_{h}^{n+1}\nabla w_{h}-w_{h}\nabla\theta_{h}^{n+1}}{(\theta_{h}^{*})^{2}}\rangle=0,$
$\langle d_{\tau}^{n+1}\mathbf{u}_{h},\mathbf{v}_{h}\rangle$ $+\mathbf{c}_{skw}(\mathbf{u}_{h}^{*},\mathbf{u}_{h}^{n+1/2},\mathbf{v}_{h})+\langle\eta^{*}\mathrm{D}\mathbf{u}_{h}^{n+1/2},\mathrm{D}\mathbf{v}_{h}\rangle-\langle\pi_{h}^{n+1},\operatorname{div}(\mathbf{v}_{h})\rangle$
$+\langle\tfrac{\phi_{h}^{*}}{\theta_{h}^{n+1}}\nabla\mu_{h}^{n+1}-\boldsymbol{\sigma}_{h}^{*}\tfrac{\nabla\theta_{h}^{n+1}}{\theta_{h}^{n+1}}-(s_{h}^{*}+\phi_{h}^{*}\mu^{*}_{h})\tfrac{\nabla\theta_{h}^{n+1}}{(\theta_{h}^{*})^{2}},\mathbf{v}_{h}\rangle=0,$
00 0 $=\langle\operatorname{div}(\mathbf{u}_{h}^{n+1/2}),q_{h}\rangle$
holds for $(\psi_{h},\xi_{h},\mathbf{v}_{h},q_{h},w_{h})\in\mathcal{V}_{h}\times\mathcal{V}_{h}\times\mathcal{X}_{h}^{d}\times\mathcal{Q}_{h}\times\mathcal{V}_{h}^{+}$. Here $g^{*}$ denotes an evaluation of $g$ at any $t\in\{t^{n},t^{n+1}\}$, but all terms have to be evaluated at the same point in time.
Theorem 2.2
For any solution $(\phi_{h},\mu_{h},\mathbf{u}_{h},\pi_{h},\theta_{h})$ of Problem 1 discrete mass and total energy conservation as well as entropy production holds, i.e.
$\langle\phi^{n+1}-\phi^{0},1\rangle=0,\qquad\langle\tfrac{1}{2}|\mathbf{u}^{n+1}_{h}|^{2}+e(\phi_{h}^{n+1},\theta_{h}^{n+1})-\tfrac{1}{2}|\mathbf{u}^{0}_{h}|^{2}-e(\phi_{h}^{0},\theta_{h}^{0},1\rangle=0,$
$\langle s(\phi_{h}^{n+1},\theta_{h}^{n+1})-s(\phi_{h}^{0},\theta_{h}^{0}),1\rangle=\tau\sum_{k=0}^{n_{T}}\mathcal{D}_{\theta^{n+1}_{h},\mathbf{L}^{*}}(\mathrm{D}\mathbf{u}^{n+1}_{h},\mu_{n}^{n+1},\theta_{h}^{n+1})+\sum_{k=0}^{n_{T}}\mathcal{D}_{num}^{k},$
where the numerical dissipation satisfies $\mathcal{D}_{num}^{n}\geq 0$ and is given by
$\mathcal{D}_{num}^{n}$ $=\tfrac{\gamma}{2}\|\nabla d^{n+1}\phi_{h}\|^{2}-\partial_{\theta\theta}\Psi(\phi^{n}_{h},\xi^{3}_{h})(d^{n+1}\theta_{h})^{2}$
$+(\partial_{\phi\phi}\Psi_{vex}(\xi^{1}_{h},\theta^{n+1}_{h})-\partial_{\phi\phi}\Psi_{cav}(\xi^{2}_{h},\theta^{n+1}_{h}))(d^{n+1}\phi_{h})^{2}.$
Proof
For total energy conservation we insert $\mathbf{v}=\mathbf{u}^{n+1/2},w=1$ together with cancellation and the skew-symmetric of $\mathbf{c}_{skew}(\mathbf{u}_{h}^{*},\mathbf{u}_{h}^{n+1/2},\mathbf{u}_{h}^{n+1/2})=0$ we obtain
$\langle\mathbf{u}_{h}^{n+1}\cdot\mathbf{u}_{h}^{n+1/2}-\mathbf{u}_{h}^{n+1/2}\cdot\mathbf{u}_{h}^{n}+e_{h}^{n+1}-e_{h}^{n},1\rangle=0.$
Using the algebraic identity $a(a+b)-(a+b)b=a^{2}-b^{2}$ we obtain after rearrangement
$\langle\tfrac{1}{2}|\mathbf{u}^{n+1}_{h}|+e_{h}^{n+1}-\tfrac{1}{2}|\mathbf{u}^{n}_{h}|^{2}-e_{h}^{n},1\rangle=0.$
For the entropy production we compute
$\langle s_{h}^{n+1}-s_{h}^{n},1\rangle=$ $\langle\theta_{h}^{n+1}e_{h}^{n+1}-\theta_{h}^{n}e_{h}^{n}-\Psi_{h}^{n+1}+\Psi_{h}^{n}-\tfrac{\gamma}{2}|\nabla\phi_{h}^{n+1}|^{2}+\tfrac{\gamma}{2}|\nabla\phi_{h}^{n}|^{2},1\rangle$
$=$ $\langle d^{n+1}e_{h},\theta_{h}^{n+1}\rangle-\gamma\langle\nabla\phi_{h}^{n+1},\nabla d^{n+1}\phi_{h}\rangle-\langle\Psi_{h}^{n+1}-\Psi_{h}^{n},1\rangle$
$+\langle e_{h}^{n},d^{n+1}\theta_{h}\rangle+\tfrac{\gamma}{2}|\nabla d^{n+1}\phi_{h}|^{2}.$
Adding ±$\pm\partial_{\phi}\Psi(\phi_{h}^{n+1},\phi_{h}^{n},\theta_{h}^{n+1})d^{n+1}\phi_{h}$ and insertion of $\xi_{h}=d^{n+1}\phi_{h}$ yields
$\langle s_{h}^{n+1}-s_{h}^{n},1\rangle=$ $\langle d^{n+1}e_{h},\theta_{h}^{n+1}\rangle-\langle\mu_{h}^{n+1},d^{n+1}\phi_{h}\rangle-\langle\Psi_{h}^{n+1}-\Psi_{h}^{n}-e_{h}^{n}d^{n+1}\theta_{h},1\rangle$
$+\langle\partial_{\phi}\Psi(\phi_{h}^{n+1},\phi_{h}^{n},\theta_{h}^{n+1}),d^{n+1}\phi_{h}\rangle+\tfrac{\gamma}{2}\|\nabla d^{n+1}\phi_{h}\|_{0}^{2}$
$=$ $\langle d^{n+1}e_{h},\theta_{h}^{n+1}\rangle-\langle\mu_{h}^{n+1},d^{n+1}\phi_{h}\rangle+\mathcal{D}_{num}^{n}.$
Insert $\psi_{h}=-\tau\mu_{h}^{n+1},w_{h}=\tau\theta_{h}^{n+1}$ into the discrete formulation yields by the same calculations as in the continuous case for the entropy production equation. Finally, for the numerical dissipation we add ±$\pm\Psi_{h}(\phi_{h}^{n},\theta_{h}^{n+1})$ which yields
$\mathcal{D}_{num}^{n}$ $=\int_{\Omega}-\Psi_{h}^{n+1}+\Psi_{h}(\phi_{h}^{n},\theta_{h}^{n+1})-e_{h}^{n}d^{n+1}\theta_{h}+\tfrac{\gamma}{2}|\nabla d^{n+1}\phi_{h}|^{2}$
$-\Psi_{h}(\phi_{h}^{n},\theta_{h}^{n+1})-\Psi_{h}^{n}+\partial_{\phi}\Psi(\phi_{h}^{n+1},\phi_{h}^{n},\theta_{h}^{n+1})d^{n+1}\phi_{h}$
$=\int_{\Omega}\tfrac{\gamma}{2}\|\nabla d^{n+1}\phi_{h}^{n+1}\|^{2}-\partial_{\theta\theta}\Psi(\phi^{n}_{h},\xi^{3}_{h})(d^{n+1}\theta_{h})^{2}$
$+(\partial_{\phi\phi}\Psi_{vex}(\xi^{1}_{h},\theta^{n+1}_{h})-\partial_{\phi\phi}\Psi_{cav}(\xi^{2}_{h},\theta^{n+1}_{h}))(d^{n+1}\phi_{h}^{n+1})^{2},$
where $\xi^{1}_{h},\xi^{2}_{h}$ are convex combinations of $\phi_{h}^{n+1},\phi_{h}^{n}$ and $\xi^{3}_{h}$ is a convex combination of $\theta_{h}^{n+1},\theta_{h}^{n}$. Using the structural assumptions on the potential $\Psi$, cf. (A4), we see that $\mathcal{D}_{num}\geq 0$ follows directly.
3 Numerical Test
For the convergence test we set $\Omega=(0,1)^{2}$ which is identified with the two-torus $\mathbb{T}^{2}$. This accounts for the periodic boundary conditions. We set $g^{*}=g^{n}$ and solve the nonlinear system with a Newton method with tolerance $10^{-12}$ in the $L^{2}$-norm. We consider the initial data
$\phi_{(}x,y)$ $=0.4+0.2\sin(2\pi x)^{2}\sin(2\pi y),\qquad\theta_{0}(x,y)=1.+0.2\sin(2\pi x)\sin(2\pi y)$
$\mathbf{u}_{0}(x,y)$ $=10^{-2}(-\sin(\pi x)^{2}\sin(2\pi y),\sin(2\pi x)\sin(\pi y)^{2})$
with the set of functionals and parameters
$\tilde{\Psi}(\phi,\theta)$ $=\log(\theta)+(2\theta-1)\phi^{2}(1-\phi)^{2}+\frac{\gamma}{2}|\nabla\phi|^{2},$
$e$ $=\tfrac{1}{\theta}+2\phi^{2}(1-\phi)^{2},\qquad s=1-\log(\theta)+1\phi^{2}(1-\phi)^{2}-\tfrac{\gamma}{2}|\nabla\phi|^{2},$
$\gamma$ $=10^{-3},\qquad\eta=10^{-3}+\tfrac{1}{40}(\phi+1)^{2},\qquad\mathbf{L}=10^{-2}\cdot\mathbf{I}.$
Since no analytical solution is available, the discretization error is estimated by comparing the computed solutions $(\phi_{h,\tau},\mu_{\rho,h,\tau},\theta_{h,\tau},\mathbf{u}_{h,\tau},p_{h,\tau})$ with those computed on uniformly refined grids, here halving space and time discretisation parameter i.e. $(\phi_{h/2,\tau/2},\mu_{\rho,h/2,\tau/2},\theta_{h/2,\tau/2},\mathbf{u}_{h/2,\tau/2},p_{h/2,\tau/2})$. The error quantities for the fully-discrete scheme are given in the energy-norm, i.e.
$e_{h,\tau}$ $=\left\|\phi_{h,\tau}-\phi_{h/2,\tau/2}\right\|_{L^{\infty}(H^{1})}^{2}+\left\|\theta_{h,\tau}-\theta_{h/2,\tau/2}\right\|_{L^{\infty}(L^{2})}^{2}+\left\|\mathbf{u}_{h,\tau}-\mathbf{u}_{h/2,\tau/2}\right\|_{L^{\infty}(L^{2})}^{2}$
$\,+\left\|\mu_{h,\tau}-\mu_{h/2,\tau/2}\right\|_{L^{2}(H^{1})}^{2}+\left\|\theta_{h,\tau}-\theta_{h/2,\tau/2}\right\|_{L^{2}(H^{1})}^{2}+\left\|\mathbf{u}_{h,\tau}-\mathbf{u}_{h/2,\tau/2}\right\|_{L^{2}(H^{1})}^{2}$
as well as the separated errors, $e^{\phi}_{h,\tau},e^{\mu}_{h,\tau},e^{\nabla\theta}_{h,\tau},e^{\nabla\mathbf{u}}_{h,\tau}$, which denote the related single quantities in the above error norm. For the convergence test we choose the discretisation parameters $h_{k}=2^{-k},\tau_{k}=10^{-3}\cdot h_{k}$ for $k=0,\ldots,4$.
Table 1: Errors and experimental orders of convergence for the CHNST system.
$k$ $e_{h,\tau}$ eoc $e^{\phi}_{h,\tau}$ eoc $e^{\mu}_{h,\tau}$ eoc $e^{\nabla\theta}_{h,\tau}$ eoc $e^{\nabla\mathbf{u}}_{h,\tau}$ eoc
00 $1.42\cdot 10^{-0}$ $1.19\cdot 10^{-0}$ $1.03\cdot 10^{-1}$ $9.34\cdot 10^{-2}$ $1.03\cdot 10^{-2}$
$1$ $5.77\cdot 10^{-1}$ 1.29 $4.74\cdot 10^{-1}$ 1.32 $5.74\cdot 10^{-2}$ 0.85 $3.84\cdot 10^{-2}$ 1.37 $6.18\cdot 10^{-3}$ 0.74
$2$ $1.59\cdot 10^{-1}$ 1.86 $1.29\cdot 10^{-1}$ 1.88 $1.69\cdot 10^{-2}$ 1.76 $1.30\cdot 10^{-2}$ 1.56 $4.64\cdot 10^{-4}$ 3.74
$3$ $3.66\cdot 10^{-2}$ 2.12 $2.88\cdot 10^{-2}$ 2.16 $4.28\cdot 10^{-3}$ 1.98 $3.47\cdot 10^{-3}$ 1.90 $2.66\cdot 10^{-5}$ 4.13
We observe in Table 1 at least first order convergence, second order for squared norms, in all norms, expect for the gradients of the velocity. Indeed, due to the piecewise quadratic elements and the Crank-Nicolson type approximation of the velocity, this can be expected if the coupling terms do not pollute the convergence rate too much. In general we would however expect that only first order convergence is valid.
4 Conclusion
In this work we have derived a fully discrete finite element scheme for the non-isothermal Cahn-Hilliard-Navier-Stokes system (1) – (3). To derive the scheme we first formulate the continuous equations into a suitable variational formulation, which allowed for a streamlined Galerkin approximation in space. The time discretisation utilises the usual convex-concave splitting for the Cahn-Hilliard component and a nonlinear implicit time discretisation for the internal energy equation. In future work, we will benchmark the scheme and consider the error analysis. Future extensions to a phase-field dependent density in the spirit of AGG will also be considered.
Acknowledgements.
Support by the German Science Foundation (DFG) via SPP 2256: Variational Methods for Predicting Complex Phenomena in Engineering Structures and Materials (project BR 7093/1-2) and via TRR 146: Multiscale Simulation Methods for Soft Matter Systems (project C3) is gratefully acknowledged. The authors would like to thank D. Trautwein for careful proof-reading of the manuscript.